# Automatically Detecting Online Deceptive Patterns

### v3

@@ 1 Introduction
Deceptive patterns, also known as ‘Dark Patterns,’ are design choices that manipulate users into making unintended decisions as they interact with applications. These patterns exploit cognitive biases and psychological vulnerabilities to influence user behavior, often in ways that benefit the service provider at the user’s expense [12, 13]. The growing use of deceptive patterns negatively impacts the quality of user experiences across various online activities, such as online purchases, engaging with social media, playing video games, or simply browsing the web [47]. The widespread nature of these patterns is well documented, with notable examples from resources like Harry Brignull’s deceptive.design [^1] https://www.deceptive.design/ and Caroline Sliders’ analysis at The Pudding [^2] https://pudding.cool/2023/05/dark-patterns/.
@@
Despite the recent push toward better web experience, fueled by user awareness, media revelations, and privacy regulations [17, 2], deceptive patterns continue to pose substantial challenges [39]. As a result, users are at risk of harm, such as financial loss [61], privacy violations [8], and the exploitation of vulnerable populations, including children [69]. In response, researchers have explored different approaches to detect and classify deceptive patterns on the web. Earlier efforts included manual analysis to analyze the distribution of deceptive patterns on the Internet [47, 30, 12]. Such approaches are infeasible at scale due to the sheer volume, dynamic nature, and diversity of website interfaces. More recent efforts include heuristic- and ML-based methods [12, 30, 47, 76, 68, 6, 18, 55, 46]. However, these approaches often exhibit limited accuracy when identifying deceptive patterns in the wild, as we show later.
@@
Recognizing the limitations of current approaches, there is a pressing need for automated tools to assist web stakeholders in navigating and mitigating online deceptive patterns. Such an automated tool must perform two primary tasks: 1) accurately identify deceptive patterns and 2) precisely localize their position within the website. The ability to both identify and localize these patterns offers several benefits to web stakeholders. First, web users can be alerted to deceptive patterns on websites they visit, enabling informed decision-making. Second, regulators can leverage such tools to identify deceptive patterns at scale, facilitating enforcement and policy development. Third, developers can gain insights into potentially problematic elements within their websites, promoting more ethical design practices [70].
@@
We propose an automated deceptive pattern detection framework, AutoBot, to address these limitations. AutoBot accurately identifies and localizes the deceptive patterns from a screenshot of a webpage. It does not rely on the underlying HTML code of the webpage, which tends to be less stable than screenshots. HTML implementation and code can vary significantly across webpages and even different accesses, while screenshots and text remain more consistent [35]. AutoBot adopts a modular design, breaking down the task into two distinct modules and leveraging existing state-of-the-art models for each. Specifically, AutoBot utilizes a Vision Module, which analyzes screenshots to accurately localize UI elements, extracting essential features into a structured, text-only format (ElementMap). It feeds this ElementMap to a Language Module that employs a Large Language Model (LLM) to analyze the ElementMap and assign a deceptive pattern based on a defined taxonomy (Section 2.1).
@@
AutoBot’s design avoids the pitfalls of directly prompting Vision Large Language Model (VLLM) for end-to-end analysis. These models, as we show in Figure 1, are known to hallucinate, leading to false positives undermining their reliability [64]. Furthermore, VLLMs currently lack the capability for accurate localization of UI elements, as shown in recent works [19, 82] and in Section 4.4.2. While fine-tuning VLLMs could potentially improve performance, the substantial demand for large annotated datasets and significant computing resources renders this approach impractical [9].
@@ 2 Background and Related Works
Web deceptive patterns refer to website interface design choices that manipulate or deceive users into making decisions they might not otherwise make [12]. Examples of such patterns include hidden costs, forced continuity, misdirection in e-commerce websites, and privacy-invasive default settings in social media platforms. Here, we present a filtered taxonomy to categorize web deceptive patterns based on existing work. We also survey existing works on detecting deceptive patterns on websites.
@@ 2.1 Taxonomy for Deceptive Patterns
Brignull et al. presented the first taxonomy of deceptive patterns in 2010 [12]. Conti et al. expanded this taxonomy to include ‘malicious interface designs’ [21]. Bösch et al. [11] introduced a similar taxonomy called ‘privacy dark patterns’, which included more privacy-centric categories such as ‘Forced Registration’ and ‘Hidden Legalese Stipulations.’
More recently, Gray et al. [30] created a unified corpus to detect deceptive designs in user interfaces, building on previous taxonomies and categories. Since then, various works have further adapted the taxonomy for specific domains. For instance, Lewis et al. [40] codified deceptive patterns for mobile apps and games, while Mathur et al. [47] adapted the taxonomy to focus on deceptive patterns present in shopping websites. Work by Chen et al. [18] and Mansur et al. [46] extended the taxonomy to detect deceptive patterns in mobile and web apps. Although prior works developed various taxonomies to classify deceptive patterns, these efforts are inconsistent and often domain-specific. To address these issues, Gray et al. [31] introduced a unified ontology of deceptive patterns integrating past literature across regulatory reports and academic works.
@@ 2.1.1 Filtered Taxonomy
We filter Gray et al.’s ontology [31] focusing only on static deceptive patterns. Figure 2 shows the filtered taxonomy, comprising four high-level categories and 11 low-level sub-types. The complete taxonomy is provided in Section 11.1. This taxonomy includes several high-level categories from Gray et al. [31]. It includes the sub-types from the Brignull et al. taxonomy [12], which map to low-level patterns in Gray et al.’s ontology. These sub-types include textual descriptions of the pattern, which we use to prompt the LLMs, as shown later. We map our taxonomy to Gray et al.’s ontology [31] in Section 11.3. Finally, the taxonomy excludes a subset of the deceptive patterns that cannot be detected through screenshots, i.e., the dynamic ones.
@@ 2.2 Detecting Deceptive Patterns
Researchers developed several mechanisms to detect and measure the prevalence of online deceptive patterns [12, 30, 47, 76, 68, 6, 18, 55, 46]. Curley et al. categorized earlier manual, automated, or semi-automated detection mechanisms [24].
@@ 2.2.1 Human-based Manual Annotations
Early efforts to identify online deceptive patterns relied on manual exploration. Brignull et al. [12] manually explored the web to compile a “Hall of Shame”: a list of websites with deceptive patterns. Gray et al. [30] expanded this corpus by performing keyword searches on social media for posts highlighting websites with deceptive patterns, which were then manually validated. While highly accurate, this methodology is limited to domain experts and lacks scalability due to manual validation requirements. To expedite the manual exploration process, Mathur et al. [47] proposed a clustering-based pipeline to group similar websites based on their text content, which is then manually inspected. Although this process accelerates data collection, it still lacks scalability.
@@ 2.2.2 Probabilistic Text & Image Models
Attempts to automate the manual inspection process include Tung et al.’s Naive Bayes classifier [76], Soe et al.’s gradient-boosted tree to flag texts showcasing deceptive patterns [68], and Adorna et al.’s combined Naive Bayes classifier and VGG-19 model to identify deceptive designs in cookie notices [6]. However, these works are often domain-specific and cannot readily generalize to new domains. Additionally, most works rely solely on text-based classifiers or heuristics, limiting their ability to detect visual deceptive patterns, such as those based on colors or trick wording.
@@ 2.2.3 Heuristic Based Methods
Recent works by Chen et al., Mansur et al., and Raju et al. [18, 46, 55] focus on automated detection of deceptive patterns in mobile apps. Chen et al. [18] and Mansur et al. [46] used predefined heuristics, limiting detection to simpler deceptive patterns like disguised ads. Prior works used the HTML code of websites to identify and detect deceptive patterns. For example, Raju et al. [55] employed rule-based source code analysis to detect patterns like ads, forced action, and nagging  [55]. However, due to the versatile and open nature of HTML, this task proves to be very challenging. While HTML implementation and code vary significantly across webpages, and even across different accesses of the same page, screenshots and text remain more consistent. As a result, recent approaches have moved completely towards text or screenshots of websites to detect and identify deceptive designs.
@@ 2.2.4 Classical ML Models
In the broader domain of Privacy, Safety, and Security (PSS), researchers have developed tools to help users navigate specific deceptive designs on websites. For instance, works by Khandelwal et al. [35, 36] enable users to find and adjust privacy settings and automatically disable non-essential cookies. Similarly, OptOutEasy by Kumar et al. [10] automatically finds opt-out links from privacy policies and surfaces them to users. These domain-specific approaches do not readily apply to deceptive pattern detection as they require retraining for each deceptive pattern.
@@ 2.2.5 Large Language Models (VLLMs)
Prior work have explored using LLMs to detect deceptive patterns. Sazid et al. [59] used GPT-3.5-Turbo to detect deceptive text with 92.57% accuracy, athough their method is limited to singular text lines, only 7 types of patterns, and ignores any surrounding visual or textual context. Similarly, Schäfer et al. [60] utilized GPT-4o to remove deceptive elements from synthetic HTML, reducing manipulativeness in 91% of cases. This approach, however, is constrained due to the versatile and non-standardized nature of HTML [71] and the challenge of fitting inflated, real-world website source code into an LLM’s context window.
More recently, a concurrent work with ours, Shi et al. [64], introduces DPGuard to automatically detect deceptive patterns from screenshots of mobile apps and websites by directly prompting VLLMs. We show later in Section 3 that VLLMs perform poorly in detecting deceptive patterns from screenshots and are often prone to hallucinations and false positives. Additionally, DPGuard only detects the presence of these deceptive patterns without any positional reference. Our work identifies deceptive patterns with significantly higher accuracy and extracts their positions, enabling us to show users exactly where these patterns occur.
@@ 3 System Overview
Prior works analyzing websites [35, 36] have primarily relied on HTML analysis. This approach faces significant challenges due to the dynamic nature of websites. Websites increasingly employ JavaScript frameworks that modify the Document Object Model (DOM) on the fly, rendering static HTML analysis insufficient. Furthermore, the diversity in the coding practices and obfuscation techniques provide additional challenges in HTML-based analyses. To address these challenges, AutoBot adopts a novel approach focusing on the invariant aspect of websites: the user experience. By leveraging the visual signals and associated text, AutoBot models how users perceive and interact with websites. This approach offers several advantages: (1) It is resilient to change in underlying technologies as it captures the actual rendered content. (2) It allows us to analyze the same information that the user encounters, providing a more accurate representation of the potential deceptive patterns.
@@
Recent advances in VLLMs offer a venue for analyzing screenshots and highlighting patterns with proper prompting. To this end, we empirically evaluate the effectiveness of GPT4.5 [5] and Gemini 2.5 Pro [72] in detecting web deceptive patterns. Our experiments showed that while these models can detect deceptive patterns, but they often hallucinate and give false positive answers. As shown in Figure 1, both Gemini 2.5 Pro and GPT-4.5 struggle to identify the deceptive patterns in screenshots. A recent work in this domain by Shi et al. [64], DPGuard, uses these models directly to detect, not localize, deceptive patterns. Consistent with Shi et al.’s evaluation, we find that DPGuard faces performance issues, struggling to generalize over a variety of websites (refer to Section 6), achieving only a macro score of 0.3452 in detecting deceptive patterns on websites.
In addition, VLLMs frequently struggle with the precise location of elements in an input image [19, 82, 41], and cannot be used to localize deceptive patterns out of the box. However, when given bounding box coordinates, these models can effectively reason about the spatial arrangements of objects [65].
@@ 4.1 Text Extraction
The Text Extraction step starts by performing Optical Character Recognition (OCR) on the website screenshot. We employ the Google Vision OCR API  [^3] https://cloud.google.com/vision/docs/ocr as it has high accuracy in extracting text from images of varying scales and resolutions. The API returns blocks of detected text. We concatenate these blocks based on proximity to form coherent text blocks. These blocks are a list of bounding boxes with their respective text content. Then, we retrieve the font size, font color, and background color of each bounding box, as illustrated in Figure 3. We calculate font size by subtracting the bottom y-coordinate from the top y-coordinate of the bounding box. To determine color information, we utilize the extcolors [16] package, extracting the most prominent color (background) and the second most prominent color (font).
This approach addresses a key limitation identified by Soe et al. [68] in the previous ML methods: the lack of representation of the UI richness that a user perceives, such as text placement and contrast between the font and background colors. Our approach captures these detailed text features, along with the relative location of text elements on the website, providing a comprehensive representation of the textual content experienced by users.

### v4

@@ 1 Introduction
Deceptive patterns, also known as ‘Dark Patterns,’ are design choices that manipulate users into making unintended decisions as they interact with applications. These patterns exploit cognitive biases and psychological vulnerabilities to influence user behavior, often in ways that benefit the service provider at the user’s expense [dark_pattern_site, brignull-2023]. The growing use of deceptive patterns negatively impacts the quality of user experiences across various online activities, such as online purchases, engaging with social media, playing video games, or simply browsing the web [mathur2019dark]. The widespread nature of these patterns is well documented, with notable examples from resources like Harry Brignull’s deceptive.design [^1] https://www.deceptive.design/ and Caroline Sliders’ analysis at The Pudding [^2] https://pudding.cool/2023/05/dark-patterns/.
@@
Despite the recent push toward better web experience, fueled by user awareness, media revelations, and privacy regulations [CPRA, GDPR2016], deceptive patterns continue to pose substantial challenges [leiser2023dark]. As a result, users are at risk of harm, such as financial loss [SchneiderWallace2023], privacy violations [Jika2023], and the exploitation of vulnerable populations, including children [luke2019]. In response, researchers have explored different approaches to detect and classify deceptive patterns on the web. Earlier efforts included manual analysis to analyze the distribution of deceptive patterns on the Internet [mathur2019dark, gray2018dark, dark_pattern_site]. Such approaches are infeasible at scale due to the sheer volume, dynamic nature, and diversity of website interfaces. More recent efforts include heuristic- and ML-based methods [dark_pattern_site, gray2018dark, mathur2019dark, Insite2023, soe2020circumvention, adorna2024developing, chen2023unveiling, raju2022smart, mansur2023aidui]. However, these approaches often exhibit limited accuracy when identifying deceptive patterns in the wild, as we show later.
@@
Recognizing the limitations of current approaches, there is a pressing need for automated tools to assist web stakeholders in navigating and mitigating online deceptive patterns. Such an automated tool must perform two primary tasks: 1) accurately identify deceptive patterns and 2) precisely localize their position within the website. The ability to both identify and localize these patterns offers several benefits to web stakeholders. First, web users can be alerted to deceptive patterns on websites they visit, enabling informed decision-making. Second, regulators can leverage such tools to identify deceptive patterns at scale, facilitating enforcement and policy development. Third, developers can gain insights into potentially problematic elements within their websites, promoting more ethical design practices [stover_how_2023].
@@
We propose an automated deceptive pattern detection framework, AutoBot, to address these limitations. AutoBot accurately identifies and localizes the deceptive patterns from a screenshot of a webpage. It does not rely on the underlying HTML code of the webpage, which tends to be less stable than screenshots. HTML implementation and code can vary significantly across webpages and even different accesses, while screenshots and text remain more consistent [khandelwal_prisec_2021]. AutoBot adopts a modular design, breaking down the task into two distinct modules and leveraging existing state-of-the-art models for each. Specifically, AutoBot utilizes a Vision Module, which analyzes screenshots to accurately localize UI elements, extracting essential features into a structured, text-only format (ElementMap). It feeds this ElementMap to a Language Module that employs a Large Language Model (LLM) to analyze the ElementMap and assign a deceptive pattern based on a defined taxonomy (Section 2.1).
@@
AutoBot’s design avoids the pitfalls of directly prompting Vision Large Language Model (VLLM) for end-to-end analysis. These models, as we show in Figure 1, are known to hallucinate, leading to false positives undermining their reliability [50Shades]. Furthermore, VLLMs currently lack the capability for accurate localization of UI elements, as shown in recent works [chen2023shikra, you2023ferret] and in Section 4.4.2. While fine-tuning VLLMs could potentially improve performance, the substantial demand for large annotated datasets and significant computing resources renders this approach impractical [AnthropicHaikuFineTune2024].
@@ 2 Background and Related Works
Web deceptive patterns refer to website interface design choices that manipulate or deceive users into making decisions they might not otherwise make [dark_pattern_site]. Examples of such patterns include hidden costs, forced continuity, misdirection in e-commerce websites, and privacy-invasive default settings in social media platforms. Here, we present a filtered taxonomy to categorize web deceptive patterns based on existing work. We also survey existing works on detecting deceptive patterns on websites.
@@ 2.1 Taxonomy for Deceptive Patterns
Brignull et al. presented the first taxonomy of deceptive patterns in 2010 [dark_pattern_site]. Conti et al. expanded this taxonomy to include ‘malicious interface designs’ [conti2010malicious]. Bösch et al. [bosch2016tales] introduced a similar taxonomy called ‘privacy dark patterns’, which included more privacy-centric categories such as ‘Forced Registration’ and ‘Hidden Legalese Stipulations.’
More recently, Gray et al. [gray2018dark] created a unified corpus to detect deceptive designs in user interfaces, building on previous taxonomies and categories. Since then, various works have further adapted the taxonomy for specific domains. For instance, Lewis et al. [lewis_gameful_2014] codified deceptive patterns for mobile apps and games, while Mathur et al. [mathur2019dark] adapted the taxonomy to focus on deceptive patterns present in shopping websites. Work by Chen et al. [chen2023unveiling] and Mansur et al. [mansur2023aidui] extended the taxonomy to detect deceptive patterns in mobile and web apps. Although prior works developed various taxonomies to classify deceptive patterns, these efforts are inconsistent and often domain-specific. To address these issues, Gray et al. [gray2023towards] introduced a unified ontology of deceptive patterns integrating past literature across regulatory reports and academic works.
@@ 2.1.1 Filtered Taxonomy
We filter Gray et al.’s ontology [gray2023towards] focusing only on static deceptive patterns. Figure 2 shows the filtered taxonomy, comprising four high-level categories and 11 low-level sub-types. The complete taxonomy is provided in Section 11.1. This taxonomy includes several high-level categories from Gray et al. [gray2023towards]. It includes the sub-types from the Brignull et al. taxonomy [dark_pattern_site], which map to low-level patterns in Gray et al.’s ontology. These sub-types include textual descriptions of the pattern, which we use to prompt the LLMs, as shown later. We map our taxonomy to Gray et al.’s ontology [gray2023towards] in Section 11.3. Finally, the taxonomy excludes a subset of the deceptive patterns that cannot be detected through screenshots, i.e., the dynamic ones.
@@ 2.2 Detecting Deceptive Patterns
Researchers developed several mechanisms to detect and measure the prevalence of online deceptive patterns [dark_pattern_site, gray2018dark, mathur2019dark, Insite2023, soe2020circumvention, adorna2024developing, chen2023unveiling, raju2022smart, mansur2023aidui]. Curley et al. categorized earlier manual, automated, or semi-automated detection mechanisms [curley2021design].
@@ 2.2.1 Human-based Manual Annotations
Early efforts to identify online deceptive patterns relied on manual exploration. Brignull et al. [dark_pattern_site] manually explored the web to compile a “Hall of Shame”: a list of websites with deceptive patterns. Gray et al. [gray2018dark] expanded this corpus by performing keyword searches on social media for posts highlighting websites with deceptive patterns, which were then manually validated. While highly accurate, this methodology is limited to domain experts and lacks scalability due to manual validation requirements. To expedite the manual exploration process, Mathur et al. [mathur2019dark] proposed a clustering-based pipeline to group similar websites based on their text content, which is then manually inspected. Although this process accelerates data collection, it still lacks scalability.
@@ 2.2.2 Probabilistic Text & Image Models
Attempts to automate the manual inspection process include Tung et al.’s Naive Bayes classifier [Insite2023], Soe et al.’s gradient-boosted tree to flag texts showcasing deceptive patterns [soe2020circumvention], and Adorna et al.’s combined Naive Bayes classifier and VGG-19 model to identify deceptive designs in cookie notices [adorna2024developing]. However, these works are often domain-specific and cannot readily generalize to new domains. Additionally, most works rely solely on text-based classifiers or heuristics, limiting their ability to detect visual deceptive patterns, such as those based on colors or trick wording.
@@ 2.2.3 Heuristic Based Methods
Recent works by Chen et al., Mansur et al., and Raju et al. [chen2023unveiling, mansur2023aidui, raju2022smart] focus on automated detection of deceptive patterns in mobile apps. Chen et al. [chen2023unveiling] and Mansur et al. [mansur2023aidui] used predefined heuristics, limiting detection to simpler deceptive patterns like disguised ads. Prior works used the HTML code of websites to identify and detect deceptive patterns. For example, Raju et al. [raju2022smart] employed rule-based source code analysis to detect patterns like ads, forced action, and nagging [raju2022smart]. However, due to the versatile and open nature of HTML, this task proves to be very challenging. While HTML implementation and code vary significantly across webpages, and even across different accesses of the same page, screenshots and text remain more consistent. As a result, recent approaches have moved completely towards text or screenshots of websites to detect and identify deceptive designs.
@@ 2.2.4 Classical ML Models
In the broader domain of Privacy, Safety, and Security (PSS), researchers have developed tools to help users navigate specific deceptive designs on websites. For instance, works by Khandelwal et al. [khandelwal_prisec_2021, khandelwal_cookie] enable users to find and adjust privacy settings and automatically disable non-essential cookies. Similarly, OptOutEasy by Kumar et al. [bannihatti_kumar_finding_2020] automatically finds opt-out links from privacy policies and surfaces them to users. These domain-specific approaches do not readily apply to deceptive pattern detection as they require retraining for each deceptive pattern.
@@ 2.2.5 Large Language Models (VLLMs)
Prior work have explored using LLMs to detect deceptive patterns. Sazid et al. [sazid2023automated] used GPT-3.5-Turbo to detect deceptive text with 92.57% accuracy, athough their method is limited to singular text lines, only 7 types of patterns, and ignores any surrounding visual or textual context. Similarly, Schäfer et al. [schafer2025don] utilized GPT-4o to remove deceptive elements from synthetic HTML, reducing manipulativeness in 91% of cases. This approach, however, is constrained due to the versatile and non-standardized nature of HTML [tan2025htmlrag] and the challenge of fitting inflated, real-world website source code into an LLM’s context window.
More recently, a concurrent work with ours, Shi et al. [50Shades], introduces DPGuard to automatically detect deceptive patterns from screenshots of mobile apps and websites by directly prompting VLLMs. We show later in Section 3 that VLLMs perform poorly in detecting deceptive patterns from screenshots and are often prone to hallucinations and false positives. Additionally, DPGuard only detects the presence of these deceptive patterns without any positional reference. Our work identifies deceptive patterns with significantly higher accuracy and extracts their positions, enabling us to show users exactly where these patterns occur.
@@ 3 System Overview
Prior works analyzing websites [khandelwal_prisec_2021, khandelwal_cookie] have primarily relied on HTML analysis. This approach faces significant challenges due to the dynamic nature of websites. Websites increasingly employ JavaScript frameworks that modify the Document Object Model (DOM) on the fly, rendering static HTML analysis insufficient. Furthermore, the diversity in the coding practices and obfuscation techniques provide additional challenges in HTML-based analyses. To address these challenges, AutoBot adopts a novel approach focusing on the invariant aspect of websites: the user experience. By leveraging the visual signals and associated text, AutoBot models how users perceive and interact with websites. This approach offers several advantages: (1) It is resilient to change in underlying technologies as it captures the actual rendered content. (2) It allows us to analyze the same information that the user encounters, providing a more accurate representation of the potential deceptive patterns.
@@
Recent advances in VLLMs offer a venue for analyzing screenshots and highlighting patterns with proper prompting. To this end, we empirically evaluate the effectiveness of GPT4.5 [achiam2023gpt] and Gemini 2.5 Pro [GoogleGemini] in detecting web deceptive patterns. Our experiments showed that while these models can detect deceptive patterns, but they often hallucinate and give false positive answers. As shown in Figure 1, both Gemini 2.5 Pro and GPT-4.5 struggle to identify the deceptive patterns in screenshots. A recent work in this domain by Shi et al. [50Shades], DPGuard, uses these models directly to detect, not localize, deceptive patterns. Consistent with Shi et al.’s evaluation, we find that DPGuard faces performance issues, struggling to generalize over a variety of websites (refer to Section 6), achieving only a macro score of 0.3452 in detecting deceptive patterns on websites.
In addition, VLLMs frequently struggle with the precise location of elements in an input image [chen2023shikra, you2023ferret, li2025towards], and cannot be used to localize deceptive patterns out of the box. However, when given bounding box coordinates, these models can effectively reason about the spatial arrangements of objects [shiri2024empirical].
@@ 4.1 Text Extraction
The Text Extraction step starts by performing Optical Character Recognition (OCR) on the website screenshot. We employ the Google Vision OCR API  [^3] https://cloud.google.com/vision/docs/ocr as it has high accuracy in extracting text from images of varying scales and resolutions. The API returns blocks of detected text. We concatenate these blocks based on proximity to form coherent text blocks. These blocks are a list of bounding boxes with their respective text content. Then, we retrieve the font size, font color, and background color of each bounding box, as illustrated in Figure 3. We calculate font size by subtracting the bottom y-coordinate from the top y-coordinate of the bounding box. To determine color information, we utilize the extcolors [extcolors] package, extracting the most prominent color (background) and the second most prominent color (font).
This approach addresses a key limitation identified by Soe et al. [soe2020circumvention] in the previous ML methods: the lack of representation of the UI richness that a user perceives, such as text placement and contrast between the font and background colors. Our approach captures these detailed text features, along with the relative location of text elements on the website, providing a comprehensive representation of the textual content experienced by users.